OpenStudy (jojo12g):

What is the prime factorization of 120? A. 2^4 • 3 B. 2^3 • 5 C. 2^2 • 3 • 5 D. 2^3 • 3 • 5

OpenStudy (aveline):

120=24*5 120=6*4*5 120=6*2*2*5 120=3*2*2*2*5 This is prime factorization...were you thinking of something else?

OpenStudy (anonymous):

The answer is D.

OpenStudy (aveline):

Yeah, she just changed the answers. 2*2*2 is 2^3, so the answer is D

OpenStudy (jojo12g):

alright thank u both so much